# suppReferences

### v2

@@ 2 Discussion
In this paper, we have demonstrated a method of achieving digital computing with optical Skyrmions using structured matter. In practice, adders of arbitrary orders can be fabricated through numerous techniques including laser polymerization of liquid crystals [38, 39, 40, 41, 42, 43], metasurfaces [44, 45, 46, 47, 48, 49, 50, 51], gradient index systems [52, 9], direct laser writing of birefringent structures in silica [53, 54], inkjet printing of liquid crystal droplets [55, 56] and more, all of which suggest the possibility of manufacturing compact microscale devices containing such adders.
As mentioned earlier, a crucial property of any such adder is that its effect on Skyrmion fields depends only on the structure of its boundary, a reflection of the topological nature of the optical Skyrmion. From an engineering perspective, this provides clear advantages in manufacturing as the tolerable margins of error in fabrication are greatly relaxed, with no further restrictions on the medium besides the continuity of its material properties. In particular, the function of the adder is independent of perturbations to material properties away from the boundary, including spatially varying anisotropic absorptions with tolerance up to a certain diattenuation and complex spatially varying retardance such as those resulting from birefringence (see [26] for the technical details). We demonstrate the robustness arising from this “topological duality” between field and matter in Extended Fig. 1, which shows the invariance of the effect of our proposed adder on the Skyrmion number of optical fields to disorder in the medium.
@@
Firstly, while our proposed medium theoretically supports wavelength division multiplexing, the generally wavelength-dependent nature of retarders may limit the number of independent operations that can be carried out simultaneously. Nevertheless, with the improving quality of on-chip comb lasers [57, 58, 59] and broadband retarders implemented through metamaterials [44, 45, 46, 47], such limitations are perhaps more a matter of engineering.
@@
Secondly, understanding the non-trivial behavior of polarization fields in waveguides is of key importance for enabling small integrated devices to make use of Skyrmions. In particular, apart from fabricating the adders themselves, it is also necessary to relay polarization information between adders for any meaningful computation to be done. For a waveguide to support the propagation of complex structured fields, it must necessarily be large enough to carry multiple modes. Moreover, work will need to be done to establish the range of Skyrmion numbers that can be achieved by an arbitrary superposition of propagating modes, and the conservation of Skyrmion number in propagation. Here, we provide a heuristic argument suggesting the feasibility of topological protection. Based on the ellipticity of the Helmholtz equation, one has, in general, that the electric field develops continuously within the waveguide. This then naturally descends onto a homotopy of Stokes fields provided there are no zeros of the field. Therefore, at least for some finite distance, we expect the Skyrmion number to remain unchanged.
@@
Thirdly, besides considerations relating to the medium, any system exploiting Skyrmions will also require auxiliary support for generating and detecting Skyrmions. On the front of generation, notice that uniform polarized light is trivially a Skyrmion of degree 0. Therefore, our proposed medium can create optical Skyrmions, where incident right circularly polarized and left circularly polarized light generate Skyrmions of degree ±$\pm k$, respectively. Regarding detection, recent developments in on-chip Stokes polarimetry [60, 61, 62] support the possibility of accurately measuring complex polarization fields on integrated circuits. However, extracting the topological number from polarimetric measurements should be done efficiently for a Skyrmion-based photonic adder to be feasible. Note that apart from evaluating the Skyrmion number integral, various other properties of the topological number can be used to determine the Skyrmion number such as counting strategies involving regular values [26]. An alternative way of detection is to exploit light-matter interactions modulated by Skyrmions, such as through optomechanical interactions between optical Skyrmions and topological solitons in liquid crystals [63].
@@
Lastly, note that depending on implementation, both passive and tuneable adders can be fabricated. The best implementation will then depend on the application, with tuneable elements providing greater flexibility at the cost of greater power consumption, possible hysteresis, and greater complexity in control.
@@
As a concluding remark, we note that with addition and subtraction possible, the remaining mathematical operations become a matter of design. However, Skyrmions also support a more natural notion of multiplication. Suppose a homogenous medium induces a mapping on the Poincaré sphere of degree $k$. The effect of such a medium on Skyrmions is equivalent to multiplication by $k$, with no other restrictions on the input field. As an example, a medium whose action on the Poincaré sphere is given by
@@
is a Skyrmion multiplier of order $k$. Lastly, note that if such a multiplier can be implemented, then division is naturally also possible, where a single rational number can be represented with a pair of Skyrmions. Division can then be done with two multiplications.

### v4

@@ 2 Discussion
In this paper, we have demonstrated a method of achieving digital computing with optical skyrmions using structured matter. In practice, adders of arbitrary orders can be fabricated through numerous techniques including laser polymerization of liquid crystals [38, 39], metasurfaces [40, 41, 42, 43, 44, 45, 46, 47], gradient index systems [9], SLM cascades [10], direct laser writing of birefringent structures in silica [48, 49], inkjet printing of liquid crystal droplets [50], compact meta-fibers [51] and more, all of which suggest the possibility of manufacturing compact microscale devices containing such adders.
As mentioned earlier, a crucial property of our proposed skyrmion photo-adder is that its effect on skyrmion fields depends only on the structure of its boundary, a reflection of the topological nature of the optical skyrmion. From an engineering perspective, this provides clear advantages in manufacturing as the tolerable margins of error in fabrication are greatly relaxed, with no further restrictions on the medium besides the continuity of its material properties. In particular, the function of the adder is independent of perturbations to material properties away from the boundary, including spatially varying anisotropic absorptions with tolerance up to a certain diattenuation and complex spatially varying retardance such as those resulting from birefringence (see Methods 1 and [26] for more details). Moreover, these conditions can be further relaxed if the framework of the generalized skyrmion number is adopted.
@@
Firstly, understanding the non-trivial behavior of polarization fields in waveguides is of key importance for enabling small integrated devices to make use of skyrmions. In particular, apart from fabricating the adders themselves, it is also necessary to relay polarization information between adders for any meaningful computation to be done. For a waveguide to support the propagation of complex structured fields, it must necessarily be large enough to carry multiple modes. Moreover, work will need to be done to establish the range of skyrmion numbers that can be achieved by an arbitrary superposition of propagating modes, and the conservation of skyrmion number in propagation. Here, we provide a heuristic argument suggesting the feasibility of topological protection. Based on the ellipticity of the Helmholtz equation, one has, in general, that the electric field develops continuously within the waveguide. This then naturally descends onto a homotopy of Stokes fields provided there are no zeros of the field. Therefore, at least for some finite distance, we expect the skyrmion number to remain unchanged.
Secondly, while our proposed medium theoretically supports wavelength division multiplexing, the generally wavelength-dependent nature of retarders may limit the number of independent operations that can be carried out simultaneously. Nevertheless, with the improving quality of on-chip comb lasers [52] and broadband retarders implemented through metamaterials [40, 41, 42, 43], such limitations are perhaps more a matter of engineering.
@@
Thirdly, besides considerations relating to the medium, any system exploiting skyrmions will also require auxiliary support for generating and detecting skyrmions. On the front of generation, notice that uniform polarized light is trivially a skyrmion of degree 0. Therefore, our proposed medium can create optical skyrmions, where incident RCP and LCP light generate skyrmions of degree ±$\pm k$, respectively. Regarding detection, recent developments in on-chip Stokes polarimetry [53] support the possibility of accurately measuring complex polarization fields on integrated circuits. However, extracting the topological number from polarimetric measurements should be done efficiently for a skyrmion-based photonic adder to be feasible. Note that apart from evaluating the skyrmion number integral, various other properties of the topological number can be used to determine the skyrmion number such as counting strategies involving regular values [26]. An alternative way of detection is to exploit light-matter interactions modulated by skyrmions, such as through optomechanical interactions between optical skyrmions and topological solitons in liquid crystals [54].
@@
Lastly, note that depending on implementation, both passive and tuneable adders can be fabricated. The best implementation will then depend on the application, with tuneable elements providing greater flexibility at the cost of greater power consumption, possible hysteresis, and greater complexity in control. We reiterate that in our work, we use SLMs to mimic a passive retarder array with disorder, rather than taking advantage of their tunability.
@@
As a concluding remark, we note that with addition and subtraction possible, the remaining mathematical operations become a matter of design. However, skyrmions also support a more natural notion of multiplication. Suppose a homogenous medium induces a mapping on the Poincaré sphere of degree $k$. The effect of such a medium on skyrmions is equivalent to multiplication by $k$, with no other restrictions on the input field. For example, a medium whose action on the Poincaré sphere is given by
@@
is a skyrmion multiplier of order $k$, while a mirror can be regarded as a multiplier of order $-1$ with the proper choice of coordinates [55]. Lastly, note that if such a multiplier can be implemented, then division is naturally also possible. One way to achieve this is by taking unique advantage of the fact that skyrmions are integer-valued and not limited to 0 and 1 like bits in conventional digital electronics. This property enables the use of more novel fundamental units of data; for example, a single rational number can be represented by either two conventional skyrmions or a generalized skyrmion $p/q\Leftrightarrow(p,q)$. Addition, multiplication and division can then be implemented by $(p_{1},q_{1})+(p_{2},q_{2})=(p_{1}q_{2}+p_{2}q_{1},q_{1}q_{2})$, $(p_{1},q_{1})\times(p_{2},q_{2})=(p_{1}p_{2},q_{1}q_{2})$ and $(p_{1},q_{1})\div(p_{2},q_{2})=(p_{1}q_{2},p_{2}q_{1})$, respectively.
Methods
1 Structured matter design using a general transformation law of skyrmions
Here, we derive a general transformation law of skyrmions from which we can observe a powerful heuristic of topological duality, namely that changes to the skyrmion number induced by a medium depend primarily on the structure of the medium on its boundary. As we will explain, this notion of duality has important implications for the design of structured matter to manipulate the skyrmion number, providing a strong tolerance to perturbations of material parameters away from the boundary.
As an aside, in this paper, we use the term structured matter as an umbrella term for the different types of matter which are used to generate and manipulate structured light. Typically, these are complex spatially varying media that can precisely module the phase (both absolute and relative), polarization and intensity of light, and can be abstractly described using Jones or Mueller calculus. Important classes of structured matter include retarders, diattenuators and depolarizers. This framework of structured matter enables the abstract design of complex components, which can then be implemented using different technologies (including metasurfaces, liquid crystal devices, and the various other examples mentioned in this paper) tailored to practical requirements such as size, resolution, tunability, reconfigurability, dynamic control, rewritability (e.g., for information storage), and more. Additionally, the topological properties of structured matter remain a largely unexplored area with potential applications in beam generation and analysis, biomedical imaging, communications, information storage, and, as demonstrated in this paper, computing.
In the following, we denote by $B_{a}(0)$ the ball of radius $a$ centered at the origin. Let $\mathcal{S},\mathcal{S}^{\prime}\colon\overline{B_{a}(0)}\longrightarrow S^{2}$ be two different continuous polarization fields that are constant on the boundary. $\mathcal{S}$ and $\mathcal{S}^{\prime}$ then descend onto the quotient $\overline{B_{a}(0)}/{\sim}\cong S^{2}$ obtained by identifying points on $\partial B_{a}(0)$, and are therefore skyrmions.
Suppose we can find a continuous homotopy $F\colon\overline{B_{a}(0)}\times[0,1]\longrightarrow S^{2}$ from $\mathcal{S}$ to $\mathcal{S}^{\prime}$. Let $S^{+}=\{(s^{1},s^{2},s^{3})\in S^{2}:s^{3}\geq 0\}$ and $S^{-}=\{(s^{1},s^{2},s^{3})\in S^{2}:s^{3}\leq 0\}$ denote the northern and southern hemispheres, respectively, and define
$\psi_{+}\colon\overline{B_{a}(0)}\longrightarrow S^{+},$ $\psi_{+}(r\cos\theta,r\sin\theta)=\left(\sqrt{1-(r/a)^{2}}\cos\theta,\sqrt{1-(r/a)^{2}}\sin\theta,\sqrt{1-(r/a)^{2}}\right)$
$\psi_{-}\colon\overline{B_{a}(0)}\longrightarrow S^{-},$ $\psi_{-}(r\cos\theta,r\sin\theta)=\left(\sqrt{1-(r/a)^{2}}\cos\theta,\sqrt{1-(r/a)^{2}}\sin\theta,-\sqrt{1-(r/a)^{2}}\right)$
so that $\psi_{+}$ is an orientation preserving diffeomorphism onto its image and $\psi_{-}$ an orientation reversing diffeomorphism onto its image. Now, let $H\colon S^{2}\times[0,1]\longrightarrow S^{2}$ be
$H(s,z)=\left\{\begin{array}[]{r l}F(\psi_{+}^{-1}(s),z),&s\in\overline{S^{+}}\\ F(a\cos\theta,a\sin\theta,z(1+s_{3})),&s\in\overline{S^{-}}\end{array}\right.,$
where $s=(s_{1},s_{2},s_{3})=(\sin\phi\cos\theta,\sin\phi\sin\theta,\cos\phi)\in S^{2}$. Then $H$ is a continuous by a gluing argument, and by the homotopy invariance of the degree, we have $\deg H(\cdot,0)=\deg H(\cdot,1).$
It is clear that $\deg H(\cdot,0)=\deg\mathcal{S}$ and $\deg H(\cdot,1)=\deg\mathcal{S}^{\prime}+\deg\partial F$ where $\partial F\colon\partial B_{a}(0)\times[0,1]/{\sim}\cong S^{2}\longrightarrow S^{2}$ with the identifications $(x,0)\sim(y,0)$ and $(x,1)\sim(y,1)$ for all $x,y\in\partial B_{a}(0)$ and $\partial F([x])=F(\iota(x))$ for the trivial inclusion $\iota\colon\partial B_{a}(0)\times[0,1]\longrightarrow\overline{B_{a}(0)}\times[0,1]$. Therefore
$\deg\mathcal{S}=\deg\mathcal{S}^{\prime}+\deg\partial F.$
If, further, $\partial F$ is $C^{1}$, we have the integral equation
$\deg\mathcal{S}=\deg\mathcal{S}^{\prime}+\int_{S^{2}}(\partial F)^{\ast}\omega$
for every normalised $\omega\in\Lambda^{2}(S^{2})$.
In particular, this proves that given a general homotopy of skyrmions, the resulting difference in skyrmion number is equal to the skyrmion number of the homotopy when restricted onto the walls of the homotoping cylinder. To make this statement more intuitive, consider the polarization field as it propagates through the medium. Initially, the boundary of the field is a single point on the Poincaré sphere. However, as the field propagates, its boundary need not remain constant valued, but instead, traces out a curve on the Poincaré sphere. Lastly this curve collapses back into a point at the output. This situation is therefore exactly analogous to the “unwrapping” of the Poincaré sphere from the north pole to the south pole that is usually used to describe the skyrmion (except here we allow for the unwrapping between two arbitrary SoPs), and indeed, the number of times the Poincaré sphere is unwrapped is exactly the difference in the skyrmion numbers of the input and output fields.
From the argument above, it is easy to understand why our proposed adder exhibits resilience to perturbations away from the boundary. This resilience arises because the change in skyrmion number depends solely on how the polarization states at the boundary are “unwrapped,” which, in turn, depends only on the material parameters at the boundary. In this sense, there is a “duality” between light and matter, where specifying the boundary condition of one imposes a corresponding condition on the other.
Another way to understand topological robustness is through the theory presented in [26], which provides precise conditions for the topological protection of optical skyrmions across a wide range of media, including spatially varying retarders, diattenuators, depolarizers, and cascades of these elements. By considering the Jones or Mueller matrix of a real adder as a composition of an ideal adder and a matrix encoding the non-idealities, the theory in [26] can be applied directly to this additional matrix.
With the procedure developed above, we may prove the following. Consider a continuous spatially varying retarder $J\colon\overline{B_{a}(0)}\longrightarrow SU(2)$ given by the parametrization
$J(\alpha,\delta,\Delta)=\begin{pmatrix}\cos^{2}(\alpha)e^{i\Delta/2}+\sin^{2}(\alpha)e^{-i\Delta/2}&2i\cos(\alpha)\sin(\alpha)\sin(\Delta/2)e^{-i\delta}\\ 2i\cos(\alpha)\sin(\alpha)\sin(\Delta/2)e^{i\delta}&\sin^{2}(\alpha)e^{i\Delta/2}+\cos^{2}(\alpha)e^{-i\Delta/2}\end{pmatrix}$
where $\mathcal{Q}\circ\alpha\colon\overline{B_{a}(0)}\longrightarrow\mathbb{R}\mathbb{P}^{1}=\mathbb{R}/(x\sim x+\pi)$ and $\mathcal{P}\circ\delta\colon\overline{B_{a}(0)}\longrightarrow S^{1}=\mathbb{R}/(x\sim x+2\pi)$ are continuous maps defining a continuously varying fast axis $(\cos(\alpha)e^{-i\delta/2},\sin(\alpha)e^{i\delta/2})$, $\mathcal{Q}$ and $\mathcal{P}$ the respective quotient maps, and $\mathcal{P}\circ\Delta\colon\overline{B_{a}(0)}\longrightarrow S^{1}$ the corresponding continuously varying retardance. Then the output field $\mathcal{S}^{\prime}$ is given by
$\mathcal{S}^{\prime}(x)=A^{T}\text{Spin}\left(J(\alpha(x),\delta(x),\Delta(x))\right)A\mathcal{S}(x)$
where $\text{Spin}\colon SU(2)\longrightarrow SO(3)$ is the usual Spin map
$\begin{pmatrix}a+bi&c+di\\ -c+di&a-bi\end{pmatrix}\mapsto\begin{pmatrix}a^{2}-b^{2}-c^{2}+d^{2}&2ab+2cd&-2ac+2bd\\ -2ab+2cd&a^{2}-b^{2}+c^{2}-d^{2}&2ad+2bc\\ 2ac+2bd&2bc-2ad&a^{2}+b^{2}-c^{2}-d^{2}\end{pmatrix}$
and
$A=\begin{pmatrix}0&1&0\\ 0&0&1\\ 1&0&0\end{pmatrix}.$
Suppose now that $\mathcal{P}\circ\delta\lvert_{\partial B_{a}(0)}=[0]$ and $\mathcal{P}\circ\Delta\lvert_{\partial B_{a}(0)}=[\pi]$. If we further restrict to inputs that satisfy $\mathcal{S}\lvert_{\partial B_{a}(0)}=(0,0,1)^{T}$, then $\mathcal{S}^{\prime}$ satisfies $\mathcal{S}^{\prime}\lvert_{\partial B_{a}(0)}=(0,0,-1)^{T}$ and is therefore a skyrmion. Moreover, we may homotope $\mathcal{S}$ to $\mathcal{S}^{\prime}$ by
$F(x,z)=A^{T}\text{Spin}\left(J(\alpha(x),\delta(x),\Delta(x)z)\right)A\mathcal{S}(x).$
One can then directly compute
$\partial F([\theta,z])=\begin{pmatrix}-\sin(2\partial\alpha(\theta))\sin((2n+1)\pi z)\\ \cos(2\partial\alpha(\theta))\sin((2n+1)\pi z)\\ \cos((2n+1)\pi z)\end{pmatrix}$
for some integer $n\in\mathbb{Z}$ and where $\partial\alpha\coloneqq\alpha\lvert_{\partial B_{a}(0)}$. Taking $\omega\in\Lambda^{2}(S^{2})$ to be the standard volume form on $S^{2}$ and assuming sufficient smoothness, we have
$\deg\partial F$ $=\int_{S^{2}}(\partial F)^{\ast}\omega$
$=\int_{0}^{2\pi}\int_{0}^{1}\frac{1}{4\pi}\partial F\cdot\left(\partial F_{\theta}\times\partial F_{z}\right)dzd\theta$
$=\int_{0}^{2\pi}\int_{0}^{1}-\frac{2n+1}{2}\sin((2n+1)\pi z)\frac{d\partial\alpha}{d\theta}dzd\theta$
[…]
$=-\deg\mathcal{Q}\circ\partial\alpha$
where $\partial F_{\theta}$ and $\partial F_{z}$ are partial derivatives of $\partial F$ with respect to $\theta$ and $z$, respectively. Therefore
$\deg\mathcal{S}^{\prime}=\deg\mathcal{S}+\deg\mathcal{Q}\circ\partial\alpha$
depends only on the number of revolutions made by the fast axis along the boundary of the medium. An identical proof can be used in the $\partial\mathcal{S}\lvert_{\partial B_{a}(0)}=(0,0,-1)^{T}$ case to show
$\deg\mathcal{S}^{\prime}=\deg\mathcal{S}-\deg\mathcal{Q}\circ\partial\alpha.$
Lastly, note that one can easily replicate this proof for more complicated geometries, in particular, the above carries directly over to any $\overline{U}$ homeomorphic to $\overline{B_{a}(0)}$.
2 Arbitrary generalized skyrmion adders
Here, we describe steps to design an arbitrary generalized skyrmion adder which performs the operation $(n)\mapsto(n+k_{1},\ldots,n+k_{j},n)$ for any collection $k_{1},\ldots,k_{j}\in\mathbb{Z}$ using elliptical retarders.
  1. 1.
    Pick an incident boundary SoP with Jones vector $J=(J_{1},J_{2})^{T}\in\mathbb{C}^{2}$.
  2. 2.
    Design a smooth curve on the Poincaré sphere which carves out $j$ components and so that the curve encircles the $i$th component $k_{i}$ times accounting for orientation.
  3. 3.
    Lift the curve from Stokes parameters to Jones vectors.
  4. 4.
    Suppose we are working with a circular domain $B_{a}(0)$. Let $(\gamma_{1},\gamma_{2})\in\mathbb{C}^{2}$ be the curve obtained in (3), and set
    $\Gamma(\theta)=\begin{pmatrix}\gamma_{1}(\theta)&-\bar{\gamma}_{2}(\theta)\\ \gamma_{2}(\theta)&\bar{\gamma}_{1}(\theta)\end{pmatrix}\begin{pmatrix}J_{1}&-\bar{J}_{2}\\ J_{2}&\bar{J}_{1}\end{pmatrix}^{\dagger}\in SU(2)$
    where represents conjugate transpose. Notice that $\Gamma(\theta)J=(\gamma_{1}(\theta),\gamma_{2}(\theta))$ by construction. Thus, we need only extend $\Gamma$ to all of $B_{a}(0)$.
  5. 5.
    Let $\Phi\colon SU(2)\longrightarrow S^{3}$ be the canonical diffeomorphism
    $\Phi\begin{pmatrix}a+bi&-c+di\\ c+di&a-bi\end{pmatrix}=\begin{pmatrix}a\\ b\\ c\\ d\end{pmatrix}.$
    We define $\tilde{\Gamma}\colon B_{a}(0)\longrightarrow SU(2)$ by
    $\Phi(\tilde{\Gamma}(r,\theta))=\frac{(r/a)\Phi(\Gamma(\theta))+(1-r/a)p}{\lVert(r/a)\Phi(\Gamma(\theta))+(1-r/a)p\rVert}$
    where $-p$ is any point not on the image of $\Phi\circ\Gamma$.
The function $\tilde{\Gamma}$ from the procedure above then describes a spatially varying Jones matrix which performs the desired additions. However, it is worth noting that the higher the order of the adder and the greater the number of numbers added simultaneously, the more challenging it becomes to realize these adders in practice. Nonetheless, we believe there is significant potential to further explore such adders and to develop adders capable of converting between generalized skyrmions.
As a side note, a generalized skyrmion can be converted back into an ordinary skyrmion using a diattenuator with fixed axes and a spatially varying extinction ratio that approaches infinity at its boundary. In this case, the resultant skyrmion number corresponds to the connected component containing the state parallel to the diattenuator’s transmissive axis. This then allows for different generalized adders to be cascaded.